How many electrons are possible in an atom with the following quantum numbers:-


n=4 and s= -1/2

Grade:11

1 Answers

ANMOL SARIN
14 years ago

Hello. You must be knowing that permissible no. of electrons ina shell is given by 2n2. For n=4 it is 32 . According to Pauli's exclusion principle half  16 electrons can have s=1/2 & remaining 16 s=-1/2. So final answer would be 16.